# AlphaApollo: Orchestrating Foundation Models and Professional Tools into a Self-Evolving System for Deep Agentic Reasoning

### v1

@@ B.1 Python Code Errors
In this section, we show cases of model-generated Python code scripts with the AlphaApollo framework. Therein, partial SyntaxError and IndentationError can be solved with the rule-based error correction component as mentioned in Section 2.2, while other errors that cannot be solved with rules are refined with the model-based error correction component. We list them as follows:
@@
    • SyntaxError(Fig. 14);
    • IndentationError (Fig. 15);
@@
    • SyntaxError (Fig. 16);
    • NameError (Fig. 17);
    • IndexError (Fig. 18);
    • TypeError (Fig. 19);
    • ValueError (Fig. 20);
    • ImportError (Fig. 21);
    • AttributeError (Fig. 22);
    • NotImplementedError (Fig. 23).
@@
B.2 Cognitive Behaviors
In this section, we provide the complete model responses for the case studies in Section 3.

### v2

@@ F.1 Python Code Errors
In this section, we show cases of model-generated Python code scripts with the AlphaApollo framework. Therein, partial SyntaxError and IndentationError can be solved with the rule-based error correction component as mentioned in Sec. B.1, while other errors that cannot be solved with rules are refined with the model-based error correction component. We list them as follows:
@@
    • SyntaxError(Figure 21);
    • IndentationError (Figure 22);
@@
    • SyntaxError (Figure 23);
    • NameError (Figure 24);
    • IndexError (Figure 25);
    • TypeError (Figure 26);
    • ValueError (Figure 27);
    • ImportError (Figure 28);
    • AttributeError (Figure 29);
    • NotImplementedError (Figure 30).
@@
F.2 Cognitive Behaviors
In this section, we provide the complete model responses in Figures 31, 32, 33, and 34 for the case studies in Sec. 3.